express each as a function of θ
a.sin(270+θ)
b.cos(π+θ)
c.tan(810+θ)
d.sin(θ-180)

Answer by lwsshak3(11628)

express each as a function of θ
a.sin(270+θ)
b.cos(π+θ)
c.tan(810+θ)
d.sin(θ-180)
..
let x=θ,pi=π
a.sin(270+x)=cosx (reference angle in quadrant 4)
b.cos(pi+x)=-cosx (reference angle in quadrant 3)
c.tan(810+x)=-cotx (reference angle in quadrant 2)
d.sin(x-180)=-sinx (reference angle in quadrant 3)